# The Social Life of Industrial Arms: How Arousal and Attention Shape Human-Robot Interaction

### v1

@@ V-A Baseline Attitudes
V-B Main Effects and Interaction Analysis
A $2\times 2$ within-subjects repeated-measures ANOVA (RM-ANOVA) was conducted on the seven dependent variables measured by the combined RoSAS–HRIES instruments, as well as on interaction durations. Assumptions of sphericity and normality were not violated. Partial eta squared ($\eta^{2}$) is reported as a measure of effect size. Significant results are summarized in Table I, and interaction effects are visualized in Fig. 3.
Attention had a significant effect on all dependent measures except discomfort and disturbance. Arousal significantly affected discomfort, sociability, and disturbance. Notably, both discomfort and sociability also exhibited significant interaction effects between arousal and attention.
V-C Post Hoc Comparisons
Bonferroni-corrected pairwise comparisons revealed the following ratings out of seven:
[…]
@@ V-C Post Hoc Comparisons
[…]
@@
V-D Interaction Time
Interaction duration (in seconds) was also analyzed. Average times by condition were:
@@ V-D Interaction Time
[…]
VI Discussion
The findings of this study highlight the importance of both arousal and attention in shaping users’ perceptions of non-anthropomorphic robotic manipulators. Significant arousal/attention interaction effects were detected for the RoSAS measure of discomfort and the HRIES measure of sociability, indicating that for low arousal conditions, the factor of attention does not appear to have a significant effect on reports of discomfort; whereas for high arousal conditions, high attention conditions significantly increase discomfort over the low attention conditions. In terms of sociability, it appears that for low arousal conditions, a robot demonstrating high attention was perceived to be more sociable than low attention, whereas, for high arousal conditions, sociability of the high attention conditions dropped, whereas reports for low attention remained relatively unchanged. Our results also indicate that the robot’s display of attention has a statistically significant effect on the RoSAS measures of warmth and competence, as well as the HRIES measures of intentionality and animacy. Variations in arousal predominantly affected feelings of disturbance.
High attention conditions yielded higher ratings of warmth, competence, and intentionality. When the robot was actively engaged with participants, its behaviours were interpreted as deliberate and controlling, suggesting a higher degree of agency. This conclusion is supported by participant feedback such as, “The robot seemed to follow my movement and then periodically look away, creating the sense that it was almost shy.” Others described it as “Cautiously curious. …Almost like a friendly dog.” These observations emphasize that when the robot “pays attention”, it is seen as more socially engaging and capable.
Increased arousal levels were associated with higher ratings of disturbance and discomfort. High arousal conditions, characterized by faster, more erratic movements, elicited feelings of unease among participants. One participant noted: “It made me feel uneasy since the tracking felt very intentional. It definitely felt like it was watching me,” while another remarked: “Clearly more aggressive. The fast-tracking made it seem hostile.” These responses indicate that while arousal can signal engagement and energy, excessive arousal may override positive social cues and trigger discomfort.
The interaction between arousal and attention further nuances our understanding of social perceptions. Under low arousal conditions, increasing attention appeared to enhance sociability and reduce discomfort. Participants in the low arousal, high attention condition described the robot as “It felt like it was tracking my head/eyes and trying to make a social connection,” suggesting that a calm demeanour combined with focused attention fosters trust and friendliness. However, when high arousal was paired with high attention, the resulting heightened energy increased discomfort and reduced sociability. One participant noted,“ This interaction made it feel as if it was an animal watching me in the wild,” while another mentioned, “It immediately felt very uncanny and creepy.” This finding suggests that the positive effects of high attention can be compromised when arousal is too intense, revealing the need for a balanced calibration.
The analysis of interaction times revealed that participants engaged for longer durations in low arousal conditions compared to high arousal conditions, suggesting that calmer robot behaviour encourages more sustained interactions. This finding aligns with the observed increase in discomfort and disturbance ratings in higher arousal conditions, potentially shortening participant engagement.
While our study provides valuable insights, several limitations must be acknowledged. First, the participant sample consisted mainly of university students within a narrow age range (18-25 years old), limiting the generalizability of the findings to a broader population. Second, our study primarily focused on short-term, freeform interactions ranging from three seconds to a few of minutes per condition. Interpretability of the findings may be limited as participants’ goals, motivations, and expectations were unconstrained. Additionally, It remains unclear how these perceptions evolve over prolonged engagement with the robot. Lastly, although the proposed system was described as platform-agnostic, all study data were collected using a UR5e manipulator; thus, participants may have attributed behaviours differently based on the form factor (industrial, mechanical arm), colouring perceptions.
Our findings suggest that the proposed attention and arousal-based system architecture effectively enables NARs to convey social and emotional states through movement and interaction. The attention engine’s ability to dynamically track and respond to human presence, combined with the arousal-modulated movement patterns, creates a foundation for socially engaging robotic behaviours. However, the results also indicate several areas for system refinement. The attention engine could be enhanced by incorporating more real-world environmental stimuli such as ambient noise levels, lighting conditions, and spatial context to better direct gaze and modulate attention, moving beyond the current virtual object-based approach. Additionally, the arousal system could be dynamically modulated based on environmental cues and interaction history, allowing the robot to transition through different emotional states that reflect its “experience” and context. These potential improvements, combined with the current findings, provide a clear direction for future iterations of the system architecture.
VII Conclusion

### v2

@@ VII-A Baseline Attitudes
VII-B Main Effects and Interaction Analysis
A $2\times 2$ within-subjects repeated-measures ANOVA revealed several significant main effects of attention. Attention significantly influenced warmth, $F(1,35)=16.97$, $p<.001$, $\eta^{2}=.327$; competence, $F(1,35)=175.35$, $p<.001$, $\eta^{2}=.834$; sociability, $F(1,35)=7.22$, $p<.01$, $\eta^{2}=.171$; animacy, $F(1,35)=18.05$, $p<.001$, $\eta^{2}=.340$; and intentionality, $F(1,35)=65.32$, $p<.001$, $\eta^{2}=.651$.
Arousal also produced significant effects, increasing discomfort, $F(1,35)=10.70$, $p<.01$, $\eta^{2}=.234$; decreasing sociability, $F(1,35)=4.75$, $p<.05$, $\eta^{2}=.120$; and increasing disturbance, $F(1,35)=7.58$, $p<.01$, $\eta^{2}=.178$.
In addition, there were significant Arousal $\times$ Attention interaction effects for discomfort, $F(1,35)=4.64$, $p<.05$, $\eta^{2}=.117$, and sociability, $F(1,35)=5.38$, $p<.05$, $\eta^{2}=.133$, indicating that the influence of arousal on these measures depended on the level of attention.
These omnibus effects were followed up with Bonferroni-corrected pairwise comparisons to examine the specific differences between conditions.
VII-C Post Hoc Comparisons
Bonferroni-corrected pairwise comparisons revealed the following effects ratings out of seven:
  • Disturbance: Higher in high ($M=3.87$, $SD=0.24$) vs. low arousal ($M=3.34$, $SD=0.22$), $p<.01$.
  • Warmth: Higher in high ($M=3.77$, $SD=0.17$) vs. low attention ($M=3.01$, $SD=0.15$), $p<.001$.
  • Competence: Higher in high ($M=5.11$, $SD=0.11$) vs. low attention ($M=3.06$, $SD=0.17$), $p<.001$.
  • Animacy: Higher in high ($M=4.10$, $SD=0.17$) vs. low attention ($M=3.44$, $SD=0.19$), $p<.001$.
  • Intentionality: Higher in high ($M=4.95$, $SD=0.13$) vs. low attention ($M=3.53$, $SD=0.16$), $p<.001$.
@@ VII-C Post Hoc Comparisons
Significant interaction effects are visualized in Fig. 3. In high attention conditions, discomfort ratings were significantly higher in high arousal ($M=3.71$, $SD=0.27$) than low arousal ($M=2.79$, $SD=0.20$), $p<.01$. Furthermore, in high attention conditions, sociability ratings were significantly lower in high arousal ($M=3.31$, $SD=0.27$) than low arousal ($M=4.17$, $SD=0.21$), $p<.01$.
@@
VII-D Interaction Time
Interaction duration (in seconds) for each trial was also analyzed. Average times by condition were:
@@ VII-D Interaction Time
Arousal level had a significant effect on interaction duration ($F=4.66$, $p<.05$, $\eta^{2}=.124$), with longer interactions observed in low arousal conditions ($M=71.96$, $SD=4.83$) than in high arousal conditions ($M=62.21$, $SD=4.99$).
VIII Discussion
While this study yielded a broad set of findings, we focus here on the three research questions posed earlier: (1) how can NARs convey emotional character through movement, (2) which parameters most influence perceptions of emotional and social engagement, and (3) how do arousal and attention interact to shape impressions of a robot’s presence and intent. We structure this discussion around three key themes: expressive attention, arousal and discomfort, and behavioural balance.
Attention had the most consistent and positive effect on perceived intentionality, warmth, and competence. Even without anthropomorphic features, dynamic orientation toward the user appeared sufficient to suggest social awareness and intent. This aligns with prior work showing that gaze-like behaviour can elicit engagement and perceptions of agency, even in non-humanoid systems [23], [13]. Participants appeared to interpret attentional cues as responsive and deliberate rather than purely mechanical. This supports existing theories of joint attention as a potent expressive tool, even in minimal or industrial robot designs. For NARs, dynamic orientation may offer a viable alternative to anthropomorphic expressivity, conveying presence without relying on form.
In contrast, arousal broadly had negative effects on social perception. As arousal increased, participants reported higher levels of discomfort and disturbance, and rated the robot as less sociable. These findings challenge the assumption that increased energy contributes positively to perceived engagement. Instead, excessive motion speed or amplitude may signal volatility, unpredictability, or threat. Prior work show that motion legibility improves trust when movements are predictable and well-scaled to context. In our study, the high-arousal conditions may have violated those expectations, especially in a form factor typically associated with precision and stability. Without anthropomorphic cues to contextualize high-energy behaviour, users may default to interpreting it as chaotic or unsafe. For NARs, this places an upper bound on expressivity through arousal alone: too much motion energy can erode social comfort, even when paired with high attention.
The most socially successful behaviours emerged when attention was high and arousal remained low. This pairing was was consistently rated as the most sociable, least disturbing, and among the most competent. Users appear to prefer systems that are attentive without being overstimulating. This reflects broader human norms around affect regulation, where calm focus is associated with trustworthiness and emotional control. These findings reinorce principles from frameworks like ELEGNT [20], which argue that expressive movement should enhance rather than distract from the interaction. For NARs, behaviour must strike a balance between clarity and intensity. Our results suggest that subtlety, rather than raw energy, is key to sustaining positive social dynamics in robotic arms.
While this study offers valuable insights into expressive behaviour design for NARs, several limitations warrant consideration. First, interactions were short-term and took place in a controlled laboratory setting with a relatively homogeneous participant pool, that is university students aged 18 to 25. These constraints limit the ecological validity and generalizability of our findings to more divers and real-world contexts. Second, the interactions were intentionally open-ended and task free to encourage spontaneous social interpretations and reduce task-induced bias. However, we recognize that task context likely shapes how users perceive arousal and attention. Third, we employed a fixed, linear arousal parameter to enable clear experimental control. While this simplification facilitated isolating the effects of arousal, it lacks adaptability. Future research could explore dynamic arousal models, such as bio=inspired or homeostatic frameworks that better reflect contextual or user-driven changes.
IX Conclusion